OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 8y = 8 x + 2y = 2

OpenStudy (kropot72):

You can find the y variable in terms of x from the second equation and then substitute for y in the first equation. Then you can solve for x. x + 2y = 2 Subtracting x from both sides gives: 2y = 2 - x Then dividing both sides by 2 gives: y = 1 - x/2 Now can you substitute this value for y into the first equation?

OpenStudy (anonymous):

i am not good at substitution :o

OpenStudy (kropot72):

Actually the method I gave won't work, the reason being the second equation in the question is simply the first equation divided by 4 on both sides. Therefore we can only find x in terms of y by rearranging the second equation as follows: x + 2y = 2 Subtracting 2y from both sides gives: x = 2 - 2y

OpenStudy (anonymous):

how would i do this equation now

OpenStudy (kropot72):

There isn't enough information in the question to enable a single solution. There are an infinite number of solutions to x = 2 - 2y. Are you sure you copied the question correctly?